OpenStudy (anonymous):

Carrots sell for $2.10 per pound, and crackers sell for $2.90 per pound. Glen bought some carrots and some crackers. The total weight was 2.3 pounds and cost $6.03.

OpenStudy (anonymous):

A. 0.5 pounds of carrots; 1.8 pounds of crackers B. 0.8 pounds of carrots; 1.5 pounds of crackers C. 1 pound of carrots; 1.3 pounds of crackers D. 1.5 pounds of carrots; 0.8 pound of crackers

OpenStudy (anonymous):

ok so do this x + y = 2.3 and 2.10 x + 2.90y = 6.03 so we can do this x = 2.3 - y and plug in x so we get 2.10(2.3 -y ) + 2.90 y = 6.03 you can finish the rest

OpenStudy (anonymous):

Thanks!